OpenStudy (anonymous):

Find the total area of a cylinder whose radius is 2 cm and height is 4 cm

OpenStudy (nurali):

A=2πrh+2πr^2 r=2 and h=4 A= 2*π*2*4+2*π*2^2 A=?

OpenStudy (anonymous):

Area of cylinder= Curved surface area + Area of two circles at the top and the bottom = ( 2*Pi*Radius*Height)+ (2* Pi*Radius*Radius)